### v1

# Spatial Context-based Self-Supervised Learning for Handwritten Text Recognition

Abstract
Handwritten Text Recognition (HTR) is a relevant problem in computer vision, and implies unique challenges owing to its inherent variability and the rich contextualization required for its interpretation. Despite the success of Self-Supervised Learning (SSL) in computer vision, its application to HTR has been rather scattered, leaving key SSL methodologies unexplored. This work focuses on one of them, namely Spatial Context-based SSL. We investigate how this family of approaches can be adapted and optimized for HTR and propose new workflows that leverage the unique features of handwritten text. Our experiments demonstrate that the methods considered lead to advancements in the state-of-the-art of SSL for HTR in a number of benchmark cases.
Keywords Self-Supervised Learning  $\cdot$ Handwritten text recognition  $\cdot$ Spatial Context
1 Introduction
Handwritten text recognition (HTR) is the research area in the field of computer vision whose objective is to transcribe the textual content of a written manuscript into a digital machine-readable format [73]. This field not only plays a key role in the current digital era of handwriting by electronic means (such as tablets) [11], but is also of paramount relevance for the preservation, indexing and dissemination of historical manuscripts that exist solely in a physical format [56].
HTR has developed considerably over the last decade owing to the emergence of Deep Learning [57], which has greatly increased its performance. However, in order to attain competitive results, these solutions usually require large volumes of manually-labelled data, which is the principal bottleneck of this method. One means by which to alleviate this problem, Self-Supervised Learning (SSL), has recently gained considerable attention from the research community [61]. SSL employs what is termed as a pretext task to leverage collections of unlabelled data for the training of neural models in order to obtain descriptive and intelligible representations [8], thus reducing the need for large amounts of labelled data [4].
The pretext tasks can be framed in different categories according to their working principle [34, 61], with the following being some of the main existing families: (i) image generation strategies [63, 46], which focus on recovering the original distribution of the data from defined distortions or corruptions; (ii) contrastive learning methods [60, 33], whose objective is to learn representative and discernible codifications of the data, and (iii) spatial context methods [27, 58], which focus on either estimating geometric transformations performed on the data [27]i.e. spatial information—or inferring the disposition of a set of patches in which the image is divided as a puzzle [21, 58]i.e. structural information.
Nevertheless, the development of SSL in computer vision has predominantly focused on image classification [45], while other domains such as HTR have garnered noticeably less attention [36]. In this respect, while there are several research efforts that explore ad-hoc adaptations from strategies used in classification tasks [3, 81, 50] or new methods particularly devised for these cases [49, 66], there is a remarkable scarcity of proposals, thus hindering the development of the HTR field.
(a) Classic spatial context-based methods
(b) HTR spatial context-based methods
Figure 1: (a) Existing spatial methods for SSL particularly designed for image classification characteristics. (b) We propose two new approaches adapted to the specific features of HTR: (i) “flip” as a geometric transformation; (ii) “sorting” as a puzzle transformation.
This work was motivated by the relevance of spatial information in HTR tasks [49], together with the conceptual simplicity and competitive performance of spatial context-based SSL strategies, and focuses on the adaptation of a reference set of such approaches to the HTR field. More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig. 1): (i) adapting the HTR task to the original (or classification-based) SSL method, namely input adaptation, and (ii) adapting the particular SSL method to the characteristics of HTR task, namely method adaptation. These proposals are comprehensively evaluated by considering several reference HTR corpora and are compared to the reference strategies followed in the field. The results obtained show that the proposed SSL strategies not only provide better recognition rates than exclusively relying on annotated data but also outperform the existing state-of-the-art methods for SSL pre-training of HTR.
The remainder of the work is structured as follows: Section 2 contextualizes the current work within the related literature of the field, while Section 3 introduces the strategies devised for spatial context-based SSL for HTR; Section 4 presents the experimental set-up considered, and Section 5 presents and analyses the results obtained. Finally, Section 6 provides the conclusion found after carrying out this work.
2 Related work
2.1 Handwritten Text Recognition
While HTR initially relied on Hidden Markov Models [40, 23, 68, 71], advances in Machine Learning, and specifically in Deep Learning, favoured the use of Recurrent Neural Networks (RNNs) owing to their capacity to model sequential data [28, 17, 42, 31]. In particular, the use of bidirectional Long Short-Term Memory (LSTMs) networks with Connectionist Temporal Classification (CTC) [28] has been the state of the art in HTR competitions for decades [5, 70, 69, 2]. Despite this prevalence of CTC, attention-based encoder-decoder approaches [7] have recently gained popularity owing to their competitive results [37, 41, 1, 18]. For a detailed review of this field, see the work of Michael et al. [52], which provides a comprehensive study of the various sequence-to-sequence approaches for HTR.
As has occurred in many other areas, there has been a growing interest in the use of more parallelizable architectures such as the Transformer [72] by adapting the work of [22] to the HTR field [10, 64, 67, 26, 20, 65]. HTR has benefited from this adaptation, either in isolation with an encoder-decoder [43, 54, 55, 74] or in combination with the CTC objective function [75, 19, 9]. The work of Diaz et al. [20] explores universal architectures for text recognition, concluding that a CNN backbone with a Transformer encoder, a CTC-based decoder, plus an explicit language model, is the most effective strategy to date. Regardless of this progress, however, the need for large labelled corpora as a pre-training strategy in Transformer-based models has become noticeable,[43, 20, 47, 79, 48]. This has led the SSL field to attract more interest in the HTR field owing to the usual scarcity of annotated data for particular domains.
2.2 Self-Supervised Text Recognition
SSL is currently a promising alternative for the pre-training of text recognition models, especially when addressing scenarios in which there is a scarcity of labelled data [4]. In this context, the first work that tackles this issue is SeqCLR [3], which adapts the SimCLR contrastive learning method [13] to the sequential nature of text recognition. Please note that, while this work is tested for both HTR and Scene Text Recognition (STR), subsequent SSL proposals have been specifically devised in order to address the nuances of each scenario. We now introduce some of the most relevant approaches in each case.
With regard to STR, PerSEC [44] is a competitive example of a contrastive framework based on extracting low-level stroke and high-level semantic contextual spaces simultaneously via a hierarchical learning approach. SimAN [49], on the contrary, is framed within a generative strategy based on recovering the original format of an image crop by means of another non-overlapping crop of the same datum. This work highlights the close relationship between HTR and spatial context information. Other noticeable generative approaches rely on Mask Image Modelling for the pre-training of full Transformer architectures, as occurs in [50, 66, 77].
In the case of HTR, the current state of the art is defined by ChaCo [81], which takes advantage of Momentum Contrast [29, 15] to pre-train a visual feature extractor by considering a Character Unit Cropping module, which ensures that two crops of the same HTR image have an overlapping part. The same authors also show that the pre-training of only the visual feature extractor and not the sequence one improves the state of the art in semi-supervised scenarios, and this is also supported by the work of Luo et al. [49].
2.3 Spatial context-based methods
Image data inherently encodes spatial information, which can be exploited in order to pre-train recognition models. According to this, SSL approaches based on this principle constitute rather simple yet effective pretext tasks. However, despite being largely explored in classification frameworks, they have rarely been studied in text recognition. We now present the two principal families of approaches that follow this principle, in addition to highlighting the main works related to them.
In terms of geometric transformations, the work of Gidaris et. al. [27] introduced the concept of predicting rotation alteration on image data as an SSL strategy. Its success in terms of both performance and simplicity led other works to explore this principle in greater depth in order to further exploit this premise [59, 80, 14, 24, 76].
Since the relative position among patches of an image encompasses rich spatial and contextual information, there are several pretext tasks that exploit this premise, namely jigsaw puzzle strategies. Some of the most common approaches rely on predicting the relative positions of two patches of an image [21] or recognising the order of a shuffled sequence of patches of an image [58, 12, 62, 53, 78].
Spatial context is rarely exploited in image-based text recognition tasks. Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—, suggest that, if adequately adapted, text recognition models could benefit from pre-training methods of this nature. Note that our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
3 Methodology
This section introduces the general recognition framework considered, after which the proposed spatial context-based SSL methods are presented and described. Fig. 2 depicts a graphical example of the proposed pretext tasks for HTR.
Figure 2: Pre-training pipeline for the different pretext tasks proposed in the work. An image is transformed according to the method and the subsequent layers handle the task. Note that Rotation, Flip and the Jigsaw puzzle predict a single class from the sets presented, while the Sorting method predicts a sequence.
3.1 Encoder-decoder architecture
In this paper, a state-of-the-art encoder-decoder pipeline is adopted for text recognition, as shown in Fig. 3. The encoder is responsible for extracting the visual and sequential characteristics of an image, while the decoder is responsible for transforming the representation into the output text.
Figure 3: Encoder-decoder pipeline. The encoder is composed of a CNN that extracts the visual features for a given image. After a map-to-sequence operation, an RNN extracts the sequential features in order to provide the final representation. The decoder, which may in this case be either Connectionist Temporal Classification or a Transformer Decoder, transforms the representation into the output text.
Encoder. Given an image $I\in\mathbb{R}^{c\times H\times W}$ as input, where $c$ is the input channels, $H$ is the height and $W$ is the width, the first goes through a Convolutional Neural Network (CNN) $\text{cnn}:\mathbb{R}^{c\times H\times W}\rightarrow\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}$, which is responsible for extracting visual features, where $c^{\prime}$ is the number of feature maps of size $H^{\prime}\times W^{\prime}$. A map-to-sequence operation $m:\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}\rightarrow\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}$ then reshapes the features into a sequence of frames. Finally, a Recurrent Neural Network (RNN) $\text{rnn}:\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}\rightarrow\mathbb{R}^{W^{\prime}\times T}$ is responsible for extracting sequential features, where $T$ is the output size that leads to the final representation $S=[s_{1},s_{2},...,s_{W^{\prime}}]$. The encoder $f(\cdot)$ can, therefore, be expressed as $f(\cdot)=\text{rnn}(m(\text{cnn}(\cdot)))$
Decoder. Given the representation $S$, in this work two different text decoders are used: (i) a CTC decoder that separately decodes each frame of $S$, and (ii) a Transformer Decoder (TD), which is responsible for autoregressively producing the text characters on the basis of the representation $S$ and the previously generated output.
3.2 SSL pipeline
This section presents the proposed adaptation of the main existing spatial context-based pretext tasks to HTR scenarios. Note that, for each pretext task, we present two possible means of enabling the use of these techniques in HTR scenarios: (i) adapting the data to a classification-oriented framework (input adaptation), and (ii) modifying the underlying nature of the pretext task in order to allow its used in sequential-based data (method adaptation).
3.2.1 Geometric transformations
As stated previously, one of the most relevant strategies in terms of geometric transformations is the prediction of rotations when applied to an image [27]. While a transformation of this nature can be applied directly in classification scenarios, the typology of the input in HTR data makes its direct use complicated, since: i) text images are rarely squared, signifying that most rotations alter the dimensionality of the axes, and ii) applying a rotation to such elements drastically changes the semantic meaning of the axes (characteristics and sequential order). We, therefore, now present the two possible adaptations of this technique to HTR scenarios:
  • Rotation (input adaptation). Given that images in HTR rarely depict the same width, a fixed-size crop—for simplicity, we consider that a square crop of dimensions $H\times H$—of the input data must be extracted in order to be able to use this pretext task. This pre-training process can be formally described as follows: given an image $I$, a squared crop is made, obtaining $I_{square}$, after which a rotation is uniformly chosen from the set {0º, 90º, 180º, 270º}, retrieving image $I_{r}$. Finally, the CNN $cnn(\cdot)$ will, together with a linear classification layer $l(\cdot)$, be responsible for solving the pretext task.
  • Flip (method adaptation). An alternative to the aforementioned case consists of applying a flip transformation rather than the rotation transformation. The advantage of this for HTR is that it does not modify the sequential dimension of the data. Moreover, the unique asymmetry of writing, which lacks symmetry both vertically and horizontally, requires a model to have the ability to understand letter shapes, carry out recognition and discern whether a flip has been applied. Formally, given an image $I$, a crop is made in order to give the image $I_{crop}$. This element is then transformed using a flipping process that is randomly selected from the set $\phi$, Horizontal (Hor.), Vertical (Vert.), Hor. and Vert. (Both)}. Finally, the CNN $cnn(\cdot)$, together with a linear classification layer $l(\cdot)$, predicts the transformation that has been applied.
3.2.2 Puzzle transformation
One of the most prominent methods in this family is the so-called jigsaw puzzle, which consists of rearranging a set of jumbled patches from the same image by predicting the index of the correct permutation [58]. Since this proposal was conceived for classification tasks in which images depict the same dimensions, we now present the alternatives conceived in order to enable its use in HTR images with different width dimensions.
  • Jigsaw puzzle (input adaptation). The original method is maintained by extracting a square crop from the HTR image, and this is later processed by employing the jigsaw puzzle strategy. Formally, for a given image $I$, a square crop of $H\times H$ dimensions is obtained from the image, obtaining $I_{square}$; the image is initially reshaped to obtain $I_{square}^{\prime}\in\mathbb{R}^{c\times 150\times 150}$ and then divided into 4 shuffled patches of dimensions $75\times 75$; these patches are further cropped into patches of size $64\times 64$, i.e., $P=[p_{1},p_{2},p_{3},p_{4}]$, where $p_{i}\in\mathbb{R}^{1\times 64\times 64}$; each patch is subsequently processed by the CNN feature extractor, whose results are concatenated and fed into a multi-layer perceptron that estimates the permutation performed.
  • Sorting (method adaptation). The adaptation of the method proposed in this work consists of rearranging a fixed-size non-overlapping number of patches—experimental parameter—into which the HTR image is divided. Formally, given an input image $I$ and a number of patches $M$—in this case, $M\in\{2,3,4\}$—, a list of patches $P=[p_{1},...,p_{M}]$, where $p_{i}\in\mathbb{R}^{1\times H\times\lfloor W/M\rfloor}$, is obtained. $P$ is then randomly unsorted and rejoined to form the image $I_{shuffled}$. Finally, the encoder $f(\cdot)$ predicts the correct order of the patches using an auxiliary decoder.
4 Experimental setup
This section presents the datasets and figures of merit considered in order to validate the proposals presented, along with the information regarding the neural architectures used and their implementation details.
4.1 Datasets
The proposed methods are assessed using three reference public datasets for HTR: the IAM Handwritten Database [51], the CVL database(CVL) [39], and George Washington (G.W.) [25]. Tab. 1 provides a summary of the characteristics of the corpora at a word-level. For the pre-training, the SSL methods described above use only samples of the training split. We consider the Aachen split [^1] https://www.openslr.org/56/ for the IAM set and that shown in Fischer et al. [25] for Washington. Since CVL does not contain an official validation split, we use 10% of the training data for validation. Note that, in all cases, we transform the input images to grayscale and normalise them to $H=64$, maintaining the aspect ratio.
Table 1: Overview of the corpora used in this work, depicting their characteristics (number of samples, size of the alphabet, number of writers, time period, and language).
Images Alphabet Writers Period Language
IAM  [51] 79 275 78 657 Modern Eng
G.W. [25] 4 894 82 2 Historic Eng
CVL  [39] 99 904 53 310 Modern Eng, Ger
The performance of the methods was analysed by employing two widely used text recognition metrics: the Character Error Rate (CER), defined as the normalised Levenshtein distance between ground-truth and prediction, and the Word Accuracy (WAcc).
4.2 Neural architectures
As mentioned previously, the recognition framework is based on an encoder-decoder scheme. Details of the precise neural architecture used are presented below.
  • Encoder: We considered a slightly modified version of Resnet-34 [30], in which we used fewer channels in intermediate layers and reduced the aggressiveness of the pooling operations, obtaining a similar capacity in terms of parameters as in [16, 6]. We opted not to employ the ResNet-29 model used in the aforementioned works, since this particular architecture was specifically designed for STR tasks—and, more precisely, for addressing images of a fixed size of $32\times 100$—in which the aspect ratio of the images is, unlike that which occurs with HTR, not particularly relevant.
  • Recurrent feature extractor: The sequential nature of the HTR data in question was modelled by considering two stacked bidirectional LSTMs [32] with a hidden size of $T=256$.
  • Transformer decoder: We followed the vanilla decoder implementation used in Vaswani et al. [72], fixing a single decoding layer with 8 attention heads.
With regard to the training details, a maximum number of 1000 epochs with a patience of 50 on the validation loss has been considered for both the pre-training and the downstream stages. Note that, in the latter case, this patience is increased to 200 epochs when a reduced part is used. The weights of the model are optimized using the Adam method [38], fixing the learning rate values to $3\cdot 10^{-4}$ and $10^{-4}$ for the pre-training and fine-tuning phases, respectively. Only the training data is used for the self-supervised pre-training. Data augmentation is also contemplated in both stages, the details of which are provided in the Supplementary material.
5 Results
The results of our experiments are presented below, and are divided into two main sections. In the first, we analyze the quality of the representations learned for HTR by the different methods. In the second, we compare these methods in a transfer learning scenario, considering cases with both limited and abundant data for fine-tuning.
In all cases, we compare our results with those obtained by ChaCo [81], given that this method has already been shown to represent the state of the art with respect to other proposals described in Section 2.
5.1 Quality evaluation
The evaluation of SSL methods often hinges on the quality of the representations learned for the task, without the aid of any auxiliary training. This quality is critical, as it directly represents the ability of the model to learn broad and general representations. In order to assess this aspect in the context of HTR, we freeze all representation layers after pre-training so as to then fine-tune them using all the training data. Note that this fine-tuning process is restricted to the final layer that projects the representations onto the character space. By freezing the representation layers, we ensure that the figures reporting recognition accuracy can be attributed to the quality of the representations themselves, rather than enhancements in the architecture of the model or the fine-tuning process as a whole.
Table 5.1 provides details of the quality assessment, averaging across all the datasets considered, [^2] The detailed results for each dataset can be found in the Supplementary material. setting our methods against the ChaCo benchmark. Focusing on our two primary SSL blocks—Geometric and Puzzle—it will be noted that both have some configurations that outperform those of ChaCo, thus demonstrating their utility for HTR tasks. With regard to the best-performing options, the proposed geometric approach “Flip” combined with TD emerges as the most promising alternative, achieving a CER and a WAcc of 17.3 % and 67.8 %, respectively, and clearly outperforming the current state of the art of ChaCo by a noticeable margin (31 % of CER and 53 % of WAcc).
Table 2: Results obtained in the representation quality evaluation in terms of the CER and WAcc figures of merit (%) for both the proposals presented and the reference ChaCo method [81]. The up and down arrows respectively denote whether the metric is positively or negatively valued. The figures represent the average of the individual performances of the datasets contemplated, with the best values per decoding scheme and metric being highlighted in bold type.
CTC TD
CER ($\downarrow$) WAcc ($\uparrow$) CER ($\downarrow$) WAcc ($\uparrow$)
Geometric
Rotation 27.4 45.3 17.8 67.8
Flip 24.4 52.4 17.3 66.9
[…]
Jigsaw 37.3 39.9 35.7 50.3
Sorting 54.7 024.9 57.8 27.6
ChaCo [81] 44.5 33.3 31.1 53.9
Of the results reported, it is worth noting a rather counterintuitive phenomenon: the only method that leverages the sequential nature of the whole word image (“Sorting”) leads to noticeably poorer representations. This outcome is surprising, because one would expect this pretext task to provide a comprehensive view of the input, closely mirroring what the network experiences during the actual HTR task. In contrast, partial images (crops) yield more informative representations for HTR, at least from the point of view of pure pre-training. This reinforces some of the conclusions drawn in literature, which suggest that focusing on local patterns such as single characters or strokes favours the learning process [49, 81]
Concentrating on more general trends, note that the TD setting systematically makes better use of the representations learned through SSL. This trend is evident, as all the TD settings outperform their counterparts that employ CTC. The representations learned would, therefore, appear to reflect an intrinsic compatibility with the TD approach for the decoding of character sequences. Furthermore, the order of methods according to their performance remains consistent regardless of the decoder utilized, whether CTC or TD. This consistency suggests that the quality of the representations is inherently sound across different decoders, a valuable trait for the development of robust HTR systems.
5.2 Transfer learning
We now explore a transfer learning scenario in depth in order to evaluate the adaptability and efficiency of the pre-trained SSL methods. The essence of this experiment is to observe how well these methods can leverage a certain amount of labelled data—5%, 10%, or 100% of the dataset—so as to fine-tune the entire network starting from the representations learned. This fine-tuning is comprehensive, involving all the layers, which contrasts with the previous experiment in which only the final layer was fine-tuned. We aim to discern the ability of these methods to obtain useful representations in HTR as a downstream task and, ultimately, to achieve the highest possible recognition accuracy.
Table 5.2 shows the results of this experiment, providing the average WAcc across all the datasets and the average relative improvements when compared to the baseline which is, in this context, defined as the performance without any pre-training. For the sake of compactness, we provide only the best decoder (CTC or TD) for each method. [^3] The detailed results for each dataset and decoder can be found in the Supplementary material. These metrics provide a holistic view of the performance of each method, thus allowing us to compare not only the absolute accuracy achieved but also the extent to which pre-training contributes to performance gains over the baseline.
Table 3: Average results in terms of the WAcc (%) metric across all datasets for the transfer learning scheme, considering different percentages of labelled data. Both the case in which no pre-training strategy is applied and the reference ChaCo method [81] are included for the purpose of comparison. Values shown in bold type denote the best results obtained per percentage of labelled data, whereas underlined figures highlight the configurations that outperform the ChaCo reference case. The average relative improvement when compared to the No pretrain scenario is presented in parentheses for each case.
Percentage of labeled data
5% 10% 100%
Baseline
No pretrain 15.8 24.1 72.9
ChaCo [81] 24.8 ($\uparrow$ 68.3) 32.9 ($\uparrow$ 46.1) 77.2 ($\uparrow$ 6.5)
[…]
Rotation 24.0 ($\uparrow$ 63.7) 37.7 ($\uparrow$ 78.6) 79.5 ($\uparrow$ 9.4)
Flip 26.0 ($\uparrow$ 75.3) 39.4 ($\uparrow$ 90.0) 79.5 ($\uparrow$ 9.6)
Puzzle 26.5 ($\uparrow$ 103.0) 36.6 ($\uparrow$ 71.9) 78.7 ($\uparrow$ 8.4)
Sorting 15.6 ($\uparrow$ 5.4) 25.3 ($\uparrow$ 12.6) 78.3 ($\uparrow$ 7.9)
In the most data-constrained scenario utilizing merely 5% of the labelled data for fine-tuning, our proposed SSL methods not only clearly surpass the baseline, as one might expect, but also exhibits a slight but consistent advantage over the state-of-the-art ChaCo method. The only exception is “Sorting”, which performs on par with the baseline. With a modest increase to 10 % for fine-tuning, the SSL methods continue to show marked improvements over the baseline, with even greater relative gains. At this level, the proposed methods, save again “Sorting”, begin to distance themselves further from ChaCo in terms of improvement (around 5 % of WAcc).
When all available data is employed for fine-tuning, all competing approaches (including the baseline) surpass 70 % of WAcc. Despite this competitiveness, the SSL methodologies proposed in this work are still those that perform the best, consistently outperforming both the baseline and ChaCo. This new level of performance is also attained by “Sorting”, unlike that which occurred in the previous scenarios, although still below our other proposals.
Although many of our proposals reflect robust behaviour that improves on the state of the art in most cases, the “Flip” method is positioned as the best-performing alternative in general terms. This method not only outperforms others in this semi-supervised scenario but was also superior in the previous section, which focused solely on the quality of the representations learned. Notably, “Flip” is tailored specifically to HTR, thus reinforcing the value of developing and implementing strategies that are directly aligned with the unique challenges and characteristics of the task in hand.
5.3 Consistency
A key objective of our study is to provide practitioners with specific guidelines with which to train more consistent HTR models through the use of SSL. In order to illustrate the effectiveness of the proposed methods, Fig. 4 provides a compelling visual representation that shows the percentage of instances in which any of our methods outperform the state-of-the-art ChaCo for any data scenarios and architectural frameworks (TD or CTC). These bars indicate that, with the exception of the “Sorting” approach, all our methods consistently perform better than that of ChaCo in the majority of cases. The methodologies introduced in this work could, therefore, be adopted as best practices, providing a greater probability of developing robust models for specific scenarios and/or architectures in the context of HTR. The “Flip” method with cropped patches specifically stands out, making it the go-to strategy for those seeking to achieve the most robust performance.
Figure 4: Comparison of occurrences of the proposed methods in the semi-supervised scenario in which they outperform the state-of-the-art ChaCo method [81] for any data collection and architectural framework (TD or CTC). The figures are provided as a ratio with respect to the total number of experimental scenarios considered. The dashed line indicates the chance of our methods to surpass the ChaCo strategy.
6 Conclusions
In this paper, we study spatial context-based SSL methods originally devised for classification tasks and adapt them to the HTR field. We propose two possible adaptation frameworks for this purpose: (i) adjusting the data in hand in order to use the original SSL methods (input adaptation), and (ii) adapting these SSL approaches to the sequential nature of HTR (method adaptation). The results obtained when considering three benchmark HTR datasets—IAM, CVL and George Washington—show that these proposals not only improve the recognition rates of the model when not considering any pre-training, but also consistently outperform the current state-of-the-art SSL methods for HTR. Moreover, it should be noted that one of our proposals that was specifically designed for HTR (flip) consistently obtains the overall best results in every experimental scenario. Note also that our results are aligned with those of previous SSL literature on text recognition [49], which argues that there is a close relationship between HTR and spatial context owing to stroke and font style information.
The objective of this study is to provide practical guidelines in order to take advantage of SSL techniques in the context of HTR, thus combating the problem of insufficient data in this field.
Appendix A Data augmentation
We detail the data augmentation techniques applied in our study, leveraging the imgaug augmentation package [35]. Our selection is informed by seminal contributions in the field [3, 81], leading to the adoption of the subsequent alterations:
Listing 1: Data augmentation used in the SSL stage.
iaa.Sequential([
iaa.SomeOf((1, 5), [
iaa.LinearContrast((0.5, 1.0)),
iaa.GaussianBlur((0.5, 1.5)),
iaa.Sharpen(alpha=(0.0, 0.5), lightness=(0.0, 0.5)),
iaa.PiecewiseAffine(scale=(0.02, 0.03), mode=’edge’),
iaa.PerspectiveTransform(scale=(0.01, 0.02)),
], random_order=True)]
Examples of the transformations generated by the previous code are shown in Fig. 5.
Figure 5: Examples of data augmentations.
Furthermore, drawing on established practices in the literature, we also employ a soft data augmentation strategy during the supervised stage. This includes subtle cropping, linear contrast adjustments, and Gaussian blur.
Appendix B Detailed results
Owing to the extensive experimentation carried out in this paper, this section offers a complete compilation of results for each method with respect to the specific architectures employed and the datasets used for pre-training and evaluation.
In Tab. B we report the quality results in terms of the CER and WAcc metrics.
In Tab. B we provide the semi-supervised results in terms of WAcc and Single Edit Distance (ED1) metrics. ED1 expresses word-level accuracy up to a single edit operation. Note that, the exclusion of the latter metric in the manuscript was primarily due to concerns regarding clarity. Nevertheless, in this section, we opted to include it to enhance the comprehensiveness of the results presented. It is worth highlighting that the incorporation of this metric does not deviate from the conclusions drawn in the paper; instead, it serves to reinforce and bolster them.
CER WAcc
IAM CVL G.W. IAM CVL G.W.
Geometric
Rotation 27.4 19.7 35.1 53.6 56.8 25.5
Flip 23.8 20.5 28.8 56.1 57.3 43.8
[…]
Jigsaw 52.4 28.2 31.3 31.6 46.0 42.2
Sorting 79.7 53.0 31.4 19.4 12.3 43.2
ChaCo 25.2 23.7 84.7 48.3 47.0 04.7
(a) Representation quality under CTC decoder.
CER WAcc
IAM CVL G.W. IAM CVL G.W.
Geometric
Rotation 19.0 11.9 22.6 64.6 79.2 59.8
Flip 17.4 15.3 19.3 64.7 75.1 61.1
[…]
Jigsaw 49.1 19.6 38.3 35.7 70.5 44.9
Sorting 72.1 58.8 64.0 8.6 24.7 19.8
ChaCo 12.2 11.9 69.2 71.6 80.2 10.0
(b) Representation quality under TD decoder.
(c) Representation quality. The frozen part is only the visual encoder pre-trained in the SSL method. Bold values indicate the best results under the same conditions of pre-training and test.
WAcc ED1
IAM CVL G.W. IAM CVL G.W.
5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100%
Baseline
No pretrain 26.5 42.2 70.4 07.5 14.2 69.7 13.5 16.0 78.5 42.1 64.9 88.1 17.4 25.4 85.6 26.3 44.6 92.3
ChaCo 30.0 46.3 64.6 13.1 13.2 67.4 15.7 21.2 72.6 47.7 70.2 83.9 34.6 40.7 84.0 28.7 34.9 89.6
[…]
Rotation 42.6 48.9 67.7 10.0 16.4 70.1 17.0 36.0 82.4 67.7 74.4 88.2 27.2 38.8 85.9 32.9 56.9 94.0
Flip 42.1 50.6 72.1 17.1 35.0 71.1 18.8 32.5 79.1 68.4 75.6 89.4 38.7 61.1 86.8 35.9 54.3 93.2
Jigsaw 42.4 51.4 67.7 05.3 12.8 55.4 12.9 19.5 72.9 66.8 75.2 86.7 18.1 31.2 77.3 26.4 43.9 89.6
Sorting 23.9 41.6 61.1 09.3 14.7 70.0 13.8 08.9 76.6 38.1 63.9 81.1 19.3 26.2 84.7 24.1 33.1 91.4
(d) Semi-supervised results under CTC decoder.
WAcc ED1
IAM CVL G.W. IAM CVL G.W.
5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100%
Baseline
No pretrain 10.3 50.3 74.8 05.9 14.2 68.7 06.3 02.1 71.3 21.7 63.6 87.6 07.9 25.4 85.6 09.2 02.4 82.8
ChaCo 47.3 53.8 75.1 21.2 38.0 83.4 05.9 07.0 73.2 66.6 74.0 88.2 27.1 44.9 87.9 06.6 08.2 83.1
[…]
Rotation 47.0 54.5 74.5 21.5 51.8 85.5 03.6 06.9 78.5 65.6 74.8 87.2 28.3 61.4 90.6 07.7 08.7 89.5
Flip 43.8 54.3 74.6 25.2 53.1 86.1 06.0 11.0 78.0 60.2 72.0 88.1 29.7 62.8 90.7 08.9 13.5 90.3
Jigsaw 43.5 51.7 74.3 21.5 38.1 84.8 14.6 20.2 77.1 60.5 69.1 87.1 27.8 44.6 88.5 18.9 26.9 88.7
Sorting 25.2 39.8 75.2 05.4 22.0 83.6 11.9 14.2 76.1 34.5 49.4 87.2 07.0 26.5 87.0 15.4 17.2 86.6
(e) Semi-supervised results under TD decoder.
(f) Semi-supervised results. The models pre-trained with the SSL methods are fine-tuned with 5%, 10% and 100% of the labelled data. Bold values indicate the best results under the same conditions of pre-training and test.

### v2

@@
 Handwritten Text Recognition (HTR) is the research area in the field of computer vision whose objective is to transcribe the textual content of a written manuscript into a digital machine-readable format [36].
-This field not only plays a key role in the current digital era of handwriting by electronic means (such as tablets) [11], but is also of paramount relevance for the preservation, indexing and dissemination of historical manuscripts that exist solely in a physical format [56].
+This area not only plays a key role in the current digital era of handwriting by electronic means (such as tablets) [5], but is also of paramount relevance for the preservation, indexing and dissemination of historical manuscripts that exist solely in a physical format [28].
 HTR has developed considerably over the last decade owing to the emergence of Deep Learning [29], which has greatly increased its performance.
-However, in order to attain competitive results, these solutions usually require large volumes of manually-labelled data, which is the principal bottleneck of this method.
+However, to attain competitive results, these solutions usually require large volumes of manually-labelled data, which constitutes their main bottleneck.
 One means by which to alleviate this problem, Self-Supervised Learning (SSL), has recently gained considerable attention from the research community [32].
-SSL employs what is termed as a pretext task to leverage collections of unlabelled data for the training of neural models in order to obtain descriptive and intelligible representations [8], thus reducing the need for large amounts of labelled data [4].
-The pretext tasks can be framed in different categories according to their working principle [34, 61], with the following being some of the main existing families: (i) image generation strategies [63, 46], which focus on recovering the original distribution of the data from defined distortions or corruptions; (ii) contrastive learning methods [60, 33], whose objective is to learn representative and discernible codifications of the data, and (iii) spatial context methods [27, 58], which focus on either estimating geometric transformations performed on the data [27]i.e.
+SSL employs a so-called pretext task to leverage collections of unlabelled data for the training of neural models to obtain descriptive and discriminative representations [4], thus reducing the need for large amounts of labelled data.
+The pretext tasks can be framed in different categories according to their working principle [18, 32], with the following being some of the main existing families: (i) image generation strategies [24], which focus on recovering the original distribution of the data from defined distortions or corruptions; (ii) discriminative learning methods [17], whose objective is to learn representative and discernible codifications of the data.
+In the last one, we can find spatial context methods, which focus on either estimating geometric transformations performed on the data [13]i.e.
 spatial information—or inferring the disposition of a set of patches in which the image is divided as a puzzle [11, 30]i.e.
 structural information.
-Nevertheless, the development of SSL in computer vision has predominantly focused on image classification [45], while other domains such as HTR have garnered noticeably less attention [36].
-In this respect, while there are several research efforts that explore ad-hoc adaptations from strategies used in classification tasks [3, 81, 50] or new methods particularly devised for these cases [49, 66], there is a remarkable scarcity of proposals, thus hindering the development of the HTR field.
+Nevertheless, the development of SSL in computer vision has predominantly focused on image classification, whereas other domains such as HTR have garnered noticeably less attention.
+In this respect, although there are several research efforts that explore ad-hoc adaptations from strategies used in classification tasks [1, 41] or new methods particularly devised for these cases [25, 33], there is a remarkable scarcity of proposals, thus leaving substantial SSL areas unexplored hindering the development of the HTR field.
 (a) Classic spatial context-based methods
@@
 (b) We propose two new approaches adapted to the specific features of HTR: (i) “Flip” as a geometric transformation; (ii) “Sorting” as a puzzle transformation.
-This work was motivated by the relevance of spatial information in HTR tasks [49], together with the conceptual simplicity and competitive performance of spatial context-based SSL strategies, and focuses on the adaptation of a reference set of such approaches to the HTR field.
+This work was motivated by the relevance of spatial information in HTR tasks [25] together with the fact that, to our best knowledge, it has not been previously studied, and focuses on the adaptation of a reference set of such approaches to the HTR field.
 More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig.
-1): (i) adapting the HTR task to the original (or classification-based) SSL method, namely input adaptation, and (ii) adapting the particular SSL method to the characteristics of HTR task, namely method adaptation.
+1): (i) adapting the HTR task to the original SSL method, namely input adaptation, and (ii) proposing novel spatial context-based SSL methods devised to the characteristics of the HTR task, namely method adaptation.
 These proposals are comprehensively evaluated by considering several reference HTR corpora and are compared to the reference strategies followed in the field.
-The results obtained show that the proposed SSL strategies not only provide better recognition rates than exclusively relying on annotated data but also outperform the existing state-of-the-art methods for SSL pre-training of HTR.
-The remainder of the work is structured as follows: Section 2 contextualizes the current work within the related literature of the field, while Section 3 introduces the strategies devised for spatial context-based SSL for HTR; Section 4 presents the experimental set-up considered, and Section 5 presents and analyses the results obtained.
-Finally, Section 6 provides the conclusion found after carrying out this work.
+The results obtained show that the proposed SSL strategies provide state-of-the-art recognition rates compared to reference SSL approaches with less computational requirements, while maintaining conceptual simplicity.
+The remainder of the work is structured as follows: Section 2 contextualizes the current work within the related literature of the field, while Section 3 introduces the strategies devised for spatial context-based SSL for HTR; Section 4 presents the experimental set-up considered, and Section 5 presents the results obtained; Section 6 provides an in-depth analysis and discussion of the strengths and limitations of the proposed methods; finally, Section 7 concludes the work.
 2 Related work
 2.1 Handwritten Text Recognition
-While HTR initially relied on Hidden Markov Models [40, 23, 68, 71], advances in Machine Learning, and specifically in Deep Learning, favoured the use of Recurrent Neural Networks (RNNs) owing to their capacity to model sequential data [28, 17, 42, 31].
-In particular, the use of bidirectional Long Short-Term Memory (LSTMs) networks with Connectionist Temporal Classification (CTC) [28] has been the state of the art in HTR competitions for decades [5, 70, 69, 2].
-Despite this prevalence of CTC, attention-based encoder-decoder approaches [7] have recently gained popularity owing to their competitive results [37, 41, 1, 18].
-For a detailed review of this field, see the work of Michael et al.
-[52], which provides a comprehensive study of the various sequence-to-sequence approaches for HTR.
-As has occurred in many other areas, there has been a growing interest in the use of more parallelizable architectures such as the Transformer [72] by adapting the work of [22] to the HTR field [10, 64, 67, 26, 20, 65].
-HTR has benefited from this adaptation, either in isolation with an encoder-decoder [43, 54, 55, 74] or in combination with the CTC objective function [75, 19, 9].
-The work of Diaz et al.
-[20] explores universal architectures for text recognition, concluding that a CNN backbone with a Transformer encoder, a CTC-based decoder, plus an explicit language model, is the most effective strategy to date.
-Regardless of this progress, however, the need for large labelled corpora as a pre-training strategy in Transformer-based models has become noticeable,[43, 20, 47, 79, 48].
-This has led the SSL field to attract more interest in the HTR field owing to the usual scarcity of annotated data for particular domains.
+While HTR initially relied on Hidden Markov Models [34], advances in Machine Learning favored the use of Recurrent Neural Networks (RNN) owing to their capacity to model sequential data.
+In particular, the use of bidirectional Long Short-Term Memory (LSTM) networks with Connectionist Temporal Classification (CTC) [14] has been the state of the art in HTR for years [36].
+Despite this prevalence of CTC, attention-based (Att.) autorregresive encoder-decoder approaches [3], have recently gained popularity owing to their competitive results.
+For a detailed review of HTR approaches, see the work of [27].
+As in other areas, there has been a growing interest in the use of more parallelizable architectures such as the Transformer [35] by adapting the work of [12].
+HTR has benefited from this adaptation, either in isolation with fully Transformer architectures [22] or in combination with other approaches [10].
+Regardless of this progress, the need for large labelled corpora as a pre-training strategy in Transformer-based models has become noticeable [22].
+This has led the SSL field to attract more interest in the HTR field owing to the usual scarcity of annotated data.
 2.2 Self-Supervised Text Recognition
-SSL is currently a promising alternative for the pre-training of text recognition models, especially when addressing scenarios in which there is a scarcity of labelled data [4].
-In this context, the first work that tackles this issue is SeqCLR [3], which adapts the SimCLR contrastive learning method [13] to the sequential nature of text recognition.
-Please note that, while this work is tested for both HTR and Scene Text Recognition (STR), subsequent SSL proposals have been specifically devised in order to address the nuances of each scenario.
+SSL represents a promising alternative for pre-training text recognition models, especially when addressing scenarios depicting scarcity of labelled data [2].
+In this context, the first work that tackled this issue was SeqCLR [1], which adapted the SimCLR contrastive learning method [7] to the sequential nature of text recognition.
+Note that, while this work is tested for both HTR and Scene Text Recognition (STR), subsequent SSL proposals have been specifically devised to address the nuances of each scenario.
 We now introduce some of the most relevant approaches in each case.
@@
 This work highlights the close relationship between HTR and spatial context information.
-Other noticeable generative approaches rely on Mask Image Modelling for the pre-training of full Transformer architectures, as occurs in [50, 66, 77].
-In the case of HTR, the current state of the art is defined by ChaCo [81], which takes advantage of Momentum Contrast [29, 15] to pre-train a visual feature extractor by considering a Character Unit Cropping module, which ensures that two crops of the same HTR image have an overlapping part.
-The same authors also show that the pre-training of only the visual feature extractor and not the sequence one improves the state of the art in semi-supervised scenarios, and this is also supported by the work of Luo et al.
-[49].
+Other noticeable generative approaches rely on Mask Image Modelling (MIM) for the pre-training of full Transformer architectures, as occurs in the works by [33] and [38].
+In the context of HTR, the ChaCo [41] and CMT-Co [40] methods constitute representative examples of the contrastive learning paradigm.
+Both strategies take advantage of Momentum Contrast [16] to pre-train a visual feature extractor, demonstrating that focusing solely on that part rather than the sequential can improve state-of-the-art performances in semi-supervised scenarios, which is also supported by the work of [25].
+Regarding MIM, we find Text-DIAE [33], which is based on denoising autoencoders for Transformer architectures.
 2.3 Spatial context-based methods
-Image data inherently encodes spatial information, which can be exploited in order to pre-train recognition models.
-According to this, SSL approaches based on this principle constitute rather simple yet effective pretext tasks.
-However, despite being largely explored in classification frameworks, they have rarely been studied in text recognition.
-We now present the two principal families of approaches that follow this principle, in addition to highlighting the main works related to them.
-In terms of geometric transformations, the work of Gidaris et.
-al.
-[27] introduced the concept of predicting rotation alteration on image data as an SSL strategy.
-Its success in terms of both performance and simplicity led other works to explore this principle in greater depth in order to further exploit this premise [59, 80, 14, 24, 76].
-Since the relative position among patches of an image encompasses rich spatial and contextual information, there are several pretext tasks that exploit this premise, namely jigsaw puzzle strategies.
-Some of the most common approaches rely on predicting the relative positions of two patches of an image [21] or recognising the order of a shuffled sequence of patches of an image [58, 12, 62, 53, 78].
-Spatial context is rarely exploited in image-based text recognition tasks.
-Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—, suggest that, if adequately adapted, text recognition models could benefit from pre-training methods of this nature.
-Note that our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
+Image data inherently encodes spatial information, which can be exploited to pre-train recognition models.
+We now present the two main families of SSL approaches that follow this principle highlighting the most important works related to them.
+In terms of geometric transformations, the work of [13] introduced the concept of predicting rotation alteration on image data as an SSL strategy.
+Its success in terms of both performance and simplicity led other works to further explore this principle in greater depth [31, 8, 37].
+Since the relative position among patches of an image encompasses rich spatial and contextual information, there are several pretext tasks that exploit this premise, namely puzzle strategies.
+Some of the most common approaches rely on predicting the relative positions of two patches of an image [11] or recognising the order of a shuffled sequence of patches of an image [30, 6, 39].
+Despite being largely explored in classification frameworks, spatial context-based SSL is rarely exploited in image-based text recognition tasks.
+Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—suggest that, if adequately adapted, text recognition models should benefit from pre-training methods of this nature.
+Note that, our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
 3 Methodology
-This section introduces the general recognition framework considered, after which the proposed spatial context-based SSL methods are presented and described.
-Fig.
-2 depicts a graphical example of the proposed pretext tasks for HTR.
-Figure 2: Pre-training pipeline for the different pretext tasks proposed in the work.
-An image is transformed according to the method and the subsequent layers handle the task.
-Note that Rotation, Flip and the Jigsaw puzzle predict a single class from the sets presented, while the Sorting method predicts a sequence.
+In this section, we introduce the general recognition framework considered and detail the spatial context-based SSL methods proposed in this work.
 3.1 Encoder-decoder architecture
-In this paper, a state-of-the-art encoder-decoder pipeline is adopted for text recognition, as shown in Fig.
-3.
+A state-of-the-art encoder-decoder pipeline is adopted for the text recognition task, as shown in Fig.
+2.
 The encoder is responsible for extracting the visual and sequential characteristics of an image, while the decoder is responsible for transforming the representation into the output text.
@@
 The encoder is composed of a CNN that extracts the visual features for a given image.
-After a map-to-sequence operation, an RNN extracts the sequential features in order to provide the final representation.
-The decoder, which may in this case be either Connectionist Temporal Classification or a Transformer Decoder, transforms the representation into the output text.
+After a map-to-sequence operation, an RNN extracts the sequential features.
+The decoder, which may in this case be either CTC or a Transformer Decoder, transforms the representation into the output text.
 Encoder.
@@
 A map-to-sequence operation $m:\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}\rightarrow\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}$ then reshapes the features into a sequence of frames.
-Finally, a Recurrent Neural Network (RNN) $\text{rnn}:\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}\rightarrow\mathbb{R}^{W^{\prime}\times T}$ is responsible for extracting sequential features, where $T$ is the output size that leads to the final representation $S=[s_{1},s_{2},...,s_{W^{\prime}}]$.
+Finally, a RNN $\text{rnn}:\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}\rightarrow\mathbb{R}^{W^{\prime}\times T}$ is responsible for extracting sequential features, where $T$ is the output size that leads to the final representation $S=[s_{1},s_{2},...,s_{W^{\prime}}]$.
 The encoder $f(\cdot)$ can, therefore, be expressed as $f(\cdot)=\text{rnn}(m(\text{cnn}(\cdot)))$
 Decoder.
-Given the representation $S$, in this work two different text decoders are used: (i) a CTC decoder that separately decodes each frame of $S$, and (ii) a Transformer Decoder (TD), which is responsible for autoregressively producing the text characters on the basis of the representation $S$ and the previously generated output.
+Given the representation $S$, in this work two different text decoders are used: (i) a CTC decoder that separately decodes each frame of $S$, and (ii) a Transformer Decoder (TD), which is responsible for autoregressively producing the text characters from the representation S and the previously generated output.
 3.2 SSL pipeline
-This section presents the proposed adaptation of the main existing spatial context-based pretext tasks to HTR scenarios.
-Note that, for each pretext task, we present two possible means of enabling the use of these techniques in HTR scenarios: (i) adapting the data to a classification-oriented framework (input adaptation), and (ii) modifying the underlying nature of the pretext task in order to allow its used in sequential-based data (method adaptation).
+This section presents the proposed adaptations for the main existing spatial context-based pretext tasks to HTR scenarios (see Fig.
+3).
+Note that for each spatial SSL family—i.e.
+geometric and puzzle—we have chosen the most representative pretext task.
+Then, for each selected method, we present two possible means of enabling the use of these techniques in HTR scenarios: (i) adapting the data to a classification-oriented framework (input adaptation), and (ii) modifying the underlying nature of the pretext task to allow its used in HTR-based data (method adaptation).
+Figure 3: Pre-training pipeline for the different pretext tasks proposed in the work.
+An image is transformed according to the method and the subsequent layers handle the task.
+Note that “Rotation”, “Flip” and “Jigsaw” predict a single class from the sets presented, while “Sorting” predicts a sequence.
 3.2.1 Geometric transformations
 As stated previously, one of the most relevant strategies in terms of geometric transformations is the prediction of rotations when applied to an image [13].
-While a transformation of this nature can be applied directly in classification scenarios, the typology of the input in HTR data makes its direct use complicated, since: i) text images are rarely squared, signifying that most rotations alter the dimensionality of the axes, and ii) applying a rotation to such elements drastically changes the semantic meaning of the axes (characteristics and sequential order).
-We, therefore, now present the two possible adaptations of this technique to HTR scenarios:
+While a transformation of this nature can be applied directly in classification scenarios, the typology of the input in HTR data makes its direct use complicated, since: i) text images are rarely squared, signifying that most rotations alter the dimensionality of the axes; and ii) applying a rotation to such elements drastically changes the semantic meaning of the axes (characteristics and sequential order).
+Therefore, we explore two possible adaptations of this technique to HTR scenarios:
   • Rotation (input adaptation).
-Given that images in HTR rarely depict the same width, a fixed-size crop—for simplicity, we consider that a square crop of dimensions $H\times H$—of the input data must be extracted in order to be able to use this pretext task.
+Given that images in HTR rarely depict the same width, a fixed-size crop—for simplicity, we consider that a square crop of dimensions $H\times H$—of the input data must be extracted to enable the use of this pretext task.
 This pre-training process can be formally described as follows: given an image $I$, a squared crop is made, obtaining $I_{square}$, after which a rotation is uniformly chosen from the set {0º, 90º, 180º, 270º}, retrieving image $I_{r}$.
@@
 One of the most prominent methods in this family is the so-called jigsaw puzzle, which consists of rearranging a set of jumbled patches from the same image by predicting the index of the correct permutation [30].
-Since this proposal was conceived for classification tasks in which images depict the same dimensions, we now present the alternatives conceived in order to enable its use in HTR images with different width dimensions.
+Since this proposal was conceived for classification tasks in which images depict the same dimensions, we now present the proposals to enable its use in HTR images with different width dimensions.
   • Jigsaw puzzle (input adaptation).
 The original method is maintained by extracting a square crop from the HTR image, and this is later processed by employing the jigsaw puzzle strategy.
-Formally, for a given image $I$, a square crop of $H\times H$ dimensions is obtained from the image, obtaining $I_{square}$; the image is initially reshaped to obtain $I_{square}^{\prime}\in\mathbb{R}^{c\times 150\times 150}$ and then divided into 4 shuffled patches of dimensions $75\times 75$; these patches are further cropped into patches of size $64\times 64$, i.e., $P=[p_{1},p_{2},p_{3},p_{4}]$, where $p_{i}\in\mathbb{R}^{1\times 64\times 64}$; each patch is subsequently processed by the CNN feature extractor, whose results are concatenated and fed into a multi-layer perceptron that estimates the permutation performed.
+Formally, for a given image $I$, a square crop of $H\times H$ dimensions is extracted from the image, obtaining $I_{square}$; the image is initially reshaped to obtain $I_{square}^{\prime}\in\mathbb{R}^{c\times 150\times 150}$ and then divided into 4 shuffled patches of dimensions $75\times 75$; these patches are further cropped into elements of size $64\times 64$, i.e., $P=[p_{1},p_{2},p_{3},p_{4}]$, where $p_{i}\in\mathbb{R}^{1\times 64\times 64}$; each patch is subsequently processed by the CNN feature extractor, whose results are concatenated and fed into a multi-layer perceptron that estimates the permutation performed among the $24$ ($4!$) possibilities.
   • Sorting (method adaptation).
-The adaptation of the method proposed in this work consists of rearranging a fixed-size non-overlapping number of patches—experimental parameter—into which the HTR image is divided.
+The proposed method adaptation consists of rearranging a fixed-size non-overlapping number of patches—experimental parameter—into which the HTR image is divided.
 Formally, given an input image $I$ and a number of patches $M$—in this case, $M\in\{2,3,4\}$—, a list of patches $P=[p_{1},...,p_{M}]$, where $p_{i}\in\mathbb{R}^{1\times H\times\lfloor W/M\rfloor}$, is obtained.
 $P$ is then randomly unsorted and rejoined to form the image $I_{shuffled}$.
-Finally, the encoder $f(\cdot)$ predicts the correct order of the patches using an auxiliary decoder.
+Finally, the encoder $f(\cdot)$ predicts the correct index order of the patches using an auxiliary decoder.
 4 Experimental setup
-This section presents the datasets and figures of merit considered in order to validate the proposals presented, along with the information regarding the neural architectures used and their implementation details.
+This section presents the datasets and the metric considered in order to validate the proposals presented, along with the information regarding the neural architectures used and their implementation details.
 4.1 Datasets
-The proposed methods are assessed using three reference public datasets for HTR: the IAM Handwritten Database [51], the CVL database(CVL) [39], and George Washington (G.W.) [25].
-Tab.
-1 provides a summary of the characteristics of the corpora at a word-level.
+The proposed methods are assessed using three reference public datasets for HTR: the IAM Handwritten Database [26], the CVL database (CVL) [21], and RIMES [15].
+Table 1 provides a summary of the characteristics of the corpora at a word-level.
 For the pre-training, the SSL methods described above use only samples of the training split.
-We consider the Aachen split [^1] https://www.openslr.org/56/ for the IAM set and that shown in Fischer et al.
-[25] for Washington.
-Since CVL does not contain an official validation split, we use 10% of the training data for validation.
-Note that, in all cases, we transform the input images to grayscale and normalise them to $H=64$, maintaining the aspect ratio.
+We consider the Aachen split [^1] https://www.openslr.org/56/ for the IAM set.
+Since CVL does not contain an official validation split, we use 5% of the training data for validation.
+In RIMES, we ensure text-letter independence (more details in Supplementary materials).
+In all cases, we transform the input images to grayscale and normalize them to $H=64$, maintaining the aspect ratio.
 Table 1: Overview of the corpora used in this work, depicting their characteristics (number of samples, size of the alphabet, number of writers, time period, and language).
-Images Alphabet Writers Period Language
-IAM  [51] 79 275 78 657 Modern Eng
-G.W. [25] 4 894 82 2 Historic Eng
-CVL  [39] 99 904 53 310 Modern Eng, Ger
-The performance of the methods was analysed by employing two widely used text recognition metrics: the Character Error Rate (CER), defined as the normalised Levenshtein distance between ground-truth and prediction, and the Word Accuracy (WAcc).
+Images Alphabet Writers Language
+IAM 79 275 78 657 Eng
+RIMES 228 358 95 1 300 French
+CVL 99 904 53 310 Eng, Ger
+For comparative purposes with the related literature [41, 33], performance was assessed in terms of Word Accuracy (WAcc).
+This figure of merit is defined as the percentage of predicted words that exactly matches the corresponding ground truth.
 4.2 Neural architectures
-As mentioned previously, the recognition framework is based on an encoder-decoder scheme. Details of the precise neural architecture used are presented below.
-  • Encoder: We considered a slightly modified version of Resnet-34 [30], in which we used fewer channels in intermediate layers and reduced the aggressiveness of the pooling operations, obtaining a similar capacity in terms of parameters as in [16, 6].
-We opted not to employ the ResNet-29 model used in the aforementioned works, since this particular architecture was specifically designed for STR tasks—and, more precisely, for addressing images of a fixed size of $32\times 100$—in which the aspect ratio of the images is, unlike that which occurs with HTR, not particularly relevant.
-  • Recurrent feature extractor: The sequential nature of the HTR data in question was modelled by considering two stacked bidirectional LSTMs [32] with a hidden size of $T=256$.
-  • Transformer decoder: We followed the vanilla decoder implementation used in Vaswani et al. [72], fixing a single decoding layer with 8 attention heads.
+The details of the considered encoder-decoder recognition scheme are presented below:
+  • Encoder: We resorted to the ResNet-29 model used in the works of [9] and [1] to provide feasible comparison, but with slight modifications to adapt it to our input characteristics.
+More details are provided in the Supplementary material.
+  • Recurrent feature extractor: The sequential nature of the HTR data was modelled by considering two stacked bidirectional LSTMs with a hidden size of $T=256$.
+  • Transformer decoder: We followed the vanilla decoder used by [35], fixing a single decoding layer with 8 attention heads and a hidden size of 512.
 With regard to the training details, a maximum number of 1000 epochs with a patience of 50 on the validation loss has been considered for both the pre-training and the downstream stages.
-Note that, in the latter case, this patience is increased to 200 epochs when a reduced part is used.
-The weights of the model are optimized using the Adam method [38], fixing the learning rate values to $3\cdot 10^{-4}$ and $10^{-4}$ for the pre-training and fine-tuning phases, respectively.
+Note that, in the latter case, this patience is increased to 200 epochs when a reduced part of the dataset is used.
 Only the training data is used for the self-supervised pre-training.
 Data augmentation is also contemplated in both stages, the details of which are provided in the Supplementary material.
+The weights of the model are optimized using the Adam method [20], fixing the learning rate values to $3\cdot 10^{-4}$ and $10^{-4}$ for the pre-training and fine-tuning phases, respectively.
 5 Results
@@
 In the first, we analyze the quality of the representations learned for HTR by the different methods.
-In the second, we compare these methods in a transfer learning scenario, considering cases with both limited and abundant data for fine-tuning.
-In all cases, we compare our results with those obtained by ChaCo [81], given that this method has already been shown to represent the state of the art with respect to other proposals described in Section 2.
+In the second, we compare these methods in a transfer learning scenario, considering cases with both limited and abundant data for fine tuning.
+In all cases, we compare against state-of-the-art SSL methods for HTR.
 5.1 Quality evaluation
 The evaluation of SSL methods often hinges on the quality of the representations learned for the task, without the aid of any auxiliary training.
-This quality is critical, as it directly represents the ability of the model to learn broad and general representations.
-In order to assess this aspect in the context of HTR, we freeze all representation layers after pre-training so as to then fine-tune them using all the training data.
-Note that this fine-tuning process is restricted to the final layer that projects the representations onto the character space.
+This quality directly represents the ability of the model to learn broad and general representations.
+To assess this aspect in the context of HTR, we freeze all representation layers after pre-training to then fine-tune the rest of the model using all the training data.
 By freezing the representation layers, we ensure that the figures reporting recognition accuracy can be attributed to the quality of the representations themselves, rather than enhancements in the architecture of the model or the fine-tuning process as a whole.
-Table 5.1 provides details of the quality assessment, averaging across all the datasets considered, [^2] The detailed results for each dataset can be found in the Supplementary material.
-setting our methods against the ChaCo benchmark.
-Focusing on our two primary SSL blocks—Geometric and Puzzle—it will be noted that both have some configurations that outperform those of ChaCo, thus demonstrating their utility for HTR tasks.
-With regard to the best-performing options, the proposed geometric approach “Flip” combined with TD emerges as the most promising alternative, achieving a CER and a WAcc of 17.3 % and 67.8 %, respectively, and clearly outperforming the current state of the art of ChaCo by a noticeable margin (31 % of CER and 53 % of WAcc).
-Table 2: Results obtained in the representation quality evaluation in terms of the CER and WAcc figures of merit (%) for both the proposals presented and the reference ChaCo method [81].
-The up and down arrows respectively denote whether the metric is positively or negatively valued.
-The figures represent the average of the individual performances of the datasets contemplated, with the best values per decoding scheme and metric being highlighted in bold type.
-CTC TD
-CER ($\downarrow$) WAcc ($\uparrow$) CER ($\downarrow$) WAcc ($\uparrow$)
+Table 2 provides details of the quality assessment for each method and dataset.
+First, focusing on our two primary SSL blocks—Geometric and Puzzle—it will be noted that the Geometric-based methods show a clear superiority against the Puzzle-based ones while being the last still competitive in several cases, thus demonstrating the utility of spatial information in HTR tasks.
+Concerning the best-performing options, ChaCo seems to be the most effective in general terms, although the Geometric strategies are close in performance.
+In this regard, the average differences for each case between the best-performing method among the Baselines and the top Geometric method are 6.2 and 2.8 WAcc points for CTC and autoregressive decoders, respectively.
+[…]
+Method CTC Autorregresive
+IAM CVL RIMES IAM CVL RIMES
+Baselines
+SeqCLR† 39.7 66.7 63.8 51.9 74.5 79.5
+ChaCo† 70.0 75.6 79.4 72.9 77.6 84.9
+CMT-Co† 53.1 72.1 66.0 58.2 74.7 74.7
+Text-DIAE‡ - - - 71.0 78.1 -
 Geometric
-Rotation 27.4 45.3 17.8 67.8
-Flip 24.4 52.4 17.3 66.9
+Rotation‡ 65.2 54.9 85.0 67.6 77.2 82.4
+Flip‡ 65.9 55.4 82.0 50.1 67.3 57.1
+Puzzle
+Jigsaw‡ 36.4 18.1 53.9 46.2 46.5 64.7
+Sorting‡ 55.0 26.4 66.1 63.3 52.7 84.3
+5.2 Transfer learning
+We now explore a transfer learning scenario to evaluate the adaptability and efficiency of the pre-trained SSL methods.
+The essence of this experiment is to observe how well these methods can leverage a certain amount of labelled data—5%, 10%, or 100% of the dataset—to fine tune the entire network starting from the representations learned during pre-training.
+This fine-tuning process involves all layers, which contrasts with the previous experiment in which only the final layers were updated.
+We aim to discern the ability of these methods to obtain useful representations in HTR as a downstream task and, ultimately, to achieve the highest possible recognition accuracy.
+Note that, the same random samples are considered for each method when using a reduced part of a dataset.
+Table 3 presents the results of this experiment reporting the WAcc for all datasets and methods.
+For the spatial context-based SSL strategies, the Geometric approaches, particularly the “Flip" proposal, demonstrate notable superiority.
+Specifically, this method exhibits a notable advantage over state-of-the-art methods when using a reduced portion of the training data in the IAM and RIMES datasets.
+For instance, in the IAM dataset with only 10% of the data, “Flip" outperforms the previously best-performing method in each scenario by 11.4 and 11.1 points for the CTC and TD decoders, respectively.
+Although the CVL dataset does not show this superiority in data-reduced scenarios, “Flip" still surpasses the previous best method for each case when using the full dataset by 7.3 and 4 points for the CTC and TD decoders, respectively.
+These results highlight the method’s ability to achieve substantial progress in state-of-the-art performance across different scenarios.
+On the other hand, it is also worth highlighting the “Rotation" method, which, while slightly trailing “Flip" in performance, still achieves remarkable improvements over the previous state-of-the-art methods in many scenarios.
+Regarding the Puzzle-based methods, while they remain competitive, they do not demonstrate better performance over either the state-of-the-art or the Geometric-based methods.
 […]
-Jigsaw 37.3 39.9 35.7 50.3
-Sorting 54.7 024.9 57.8 27.6
-ChaCo [81] 44.5 33.3 31.1 53.9
-Of the results reported, it is worth noting a rather counterintuitive phenomenon: the only method that leverages the sequential nature of the whole word image (“Sorting”) leads to noticeably poorer representations.
-This outcome is surprising, because one would expect this pretext task to provide a comprehensive view of the input, closely mirroring what the network experiences during the actual HTR task.
+Method IAM CVL RIMES
+5 10 100 5 10 100 5 10 100
+Baselines
+SeqCLR 31.2 44.9 76.7 66.0 71.0 77.0 61.8 71.9 90.1
+ChaCo 44.3 53.1 79.8 65.6 70.3 78.1 61.3 73.0 89.2
+CMT-Co 47.7 56.0 80.1 71.9 74.5 78.2 67.1 75.4 90.1
+PerSec-CNN - - 77.9 - - 78.1 - - -
+Geometric
+Rotation 60.2 66.3 79.0 39.0 52.5 85.3 79.7 84.4 91.4
+Flip 61.7 67.4 79.1 40.6 54.6 85.5 79.8 84.2 91.1
+Puzzle
+Jigsaw 54.5 64.0 77.7 28.7 42.1 83.8 77.6 81.0 90.8
+Sorting 56.4 64.7 79.0 31.8 34.7 78.5 74.1 71.5 90.0
+(a) CTC decoder.
+Method IAM CVL RIMES
+5 10 100 5 10 100 5 10 100
+Baselines
+SeqCLR† 40.3 52.3 79.9 73.1 74.8 77.8 70.9 77.0 92.5
+ChaCo† 46.7 55.5 81.4 68.9 74.0 78.2 68.4 77.0 90.6
+CMT-Co† 50.4 55.8 81.9 73.6 76.2 78.7 68.8 77.6 91.2
+PerSec-CNN† - - 80.8 - - 80.2 - - -
+Text-DIAE‡ 49.6 58.7 80.0 47.9 68.5 87.3 - - -
+Geometric
+Rotation‡ 62.9 68.1 79.6 53.5 66.5 91.2 80.3 84.8 91.7
+Flip‡ 62.5 69.8 80.8 52.1 66.3 91.3 79.0 83.9 89.7
+Puzzle
+Jigsaw‡ 58.5 68.2 80.7 43.5 63.3 90.1 78.9 83.2 92.0
+Sorting‡ 59.2 69.2 80.8 43.4 54.2 86.2 80.4 83.9 91.7
+(b) Autorregresive decoder.
+6 Discussion
+In this section, we delve deeper into the analysis and interpretation of the results.
+For this purpose, we also visualize the feature extraction of each spatial context-based SSL method, compare efficiency of the different neural architectures, and examine the strengths and limitations of the proposed methods.
+6.1 Comparison and feature visualization
+Considering the results reported, it is worth noting a rather counterintuitive phenomenon: the only method that leverages the sequential nature of the whole word image (“Sorting”) does not lead the best representations.
+This outcome is surprising, as one would expect this pretext task to provide a comprehensive view of the input, closely mirroring what the network experiences during the actual HTR task.
 In contrast, partial images (crops) yield more informative representations for HTR, at least from the point of view of pure pre-training.
-This reinforces some of the conclusions drawn in literature, which suggest that focusing on local patterns such as single characters or strokes favours the learning process [49, 81]
-Concentrating on more general trends, note that the TD setting systematically makes better use of the representations learned through SSL.
+This argument is supported by Fig.
+4, which visualizes the feature extraction of the proposed methods as a heat map, highlighting that “Sorting" produces less effective visual features than the Geometric approaches.
+In addition, note that the “Jigsaw" method, which focuses at sub-character level (just partial views of characters), also shows a clear worse attention to the stroke patterns.
+This reinforces some of the conclusions drawn in the literature, which suggest that focusing on local patterns such as single characters or strokes favours the learning process [25, 41].
+It may be also observed that in our approaches, the TD setting systematically makes better use of the representations learned through SSL.
 This trend is evident, as all the TD settings outperform their counterparts that employ CTC.
@@
 This consistency suggests that the quality of the representations is inherently sound across different decoders, a valuable trait for the development of robust HTR systems.
-5.2 Transfer learning
-We now explore a transfer learning scenario in depth in order to evaluate the adaptability and efficiency of the pre-trained SSL methods.
-The essence of this experiment is to observe how well these methods can leverage a certain amount of labelled data—5%, 10%, or 100% of the dataset—so as to fine-tune the entire network starting from the representations learned.
-This fine-tuning is comprehensive, involving all the layers, which contrasts with the previous experiment in which only the final layer was fine-tuned.
-We aim to discern the ability of these methods to obtain useful representations in HTR as a downstream task and, ultimately, to achieve the highest possible recognition accuracy.
-Table 5.2 shows the results of this experiment, providing the average WAcc across all the datasets and the average relative improvements when compared to the baseline which is, in this context, defined as the performance without any pre-training.
-For the sake of compactness, we provide only the best decoder (CTC or TD) for each method.
-[^3] The detailed results for each dataset and decoder can be found in the Supplementary material.
-These metrics provide a holistic view of the performance of each method, thus allowing us to compare not only the absolute accuracy achieved but also the extent to which pre-training contributes to performance gains over the baseline.
-Table 3: Average results in terms of the WAcc (%) metric across all datasets for the transfer learning scheme, considering different percentages of labelled data.
-Both the case in which no pre-training strategy is applied and the reference ChaCo method [81] are included for the purpose of comparison.
-Values shown in bold type denote the best results obtained per percentage of labelled data, whereas underlined figures highlight the configurations that outperform the ChaCo reference case.
-The average relative improvement when compared to the No pretrain scenario is presented in parentheses for each case.
-Percentage of labeled data
-5% 10% 100%
-Baseline
-No pretrain 15.8 24.1 72.9
-ChaCo [81] 24.8 ($\uparrow$ 68.3) 32.9 ($\uparrow$ 46.1) 77.2 ($\uparrow$ 6.5)
-[…]
-Rotation 24.0 ($\uparrow$ 63.7) 37.7 ($\uparrow$ 78.6) 79.5 ($\uparrow$ 9.4)
-Flip 26.0 ($\uparrow$ 75.3) 39.4 ($\uparrow$ 90.0) 79.5 ($\uparrow$ 9.6)
-Puzzle 26.5 ($\uparrow$ 103.0) 36.6 ($\uparrow$ 71.9) 78.7 ($\uparrow$ 8.4)
-Sorting 15.6 ($\uparrow$ 5.4) 25.3 ($\uparrow$ 12.6) 78.3 ($\uparrow$ 7.9)
-In the most data-constrained scenario utilizing merely 5% of the labelled data for fine-tuning, our proposed SSL methods not only clearly surpass the baseline, as one might expect, but also exhibits a slight but consistent advantage over the state-of-the-art ChaCo method.
-The only exception is “Sorting”, which performs on par with the baseline.
-With a modest increase to 10 % for fine-tuning, the SSL methods continue to show marked improvements over the baseline, with even greater relative gains.
-At this level, the proposed methods, save again “Sorting”, begin to distance themselves further from ChaCo in terms of improvement (around 5 % of WAcc).
-When all available data is employed for fine-tuning, all competing approaches (including the baseline) surpass 70 % of WAcc.
-Despite this competitiveness, the SSL methodologies proposed in this work are still those that perform the best, consistently outperforming both the baseline and ChaCo.
-This new level of performance is also attained by “Sorting”, unlike that which occurred in the previous scenarios, although still below our other proposals.
-Although many of our proposals reflect robust behaviour that improves on the state of the art in most cases, the “Flip” method is positioned as the best-performing alternative in general terms.
-This method not only outperforms others in this semi-supervised scenario but was also superior in the previous section, which focused solely on the quality of the representations learned.
-Notably, “Flip” is tailored specifically to HTR, thus reinforcing the value of developing and implementing strategies that are directly aligned with the unique challenges and characteristics of the task in hand.
-5.3 Consistency
-A key objective of our study is to provide practitioners with specific guidelines with which to train more consistent HTR models through the use of SSL.
-In order to illustrate the effectiveness of the proposed methods, Fig.
-4 provides a compelling visual representation that shows the percentage of instances in which any of our methods outperform the state-of-the-art ChaCo for any data scenarios and architectural frameworks (TD or CTC).
-These bars indicate that, with the exception of the “Sorting” approach, all our methods consistently perform better than that of ChaCo in the majority of cases.
-The methodologies introduced in this work could, therefore, be adopted as best practices, providing a greater probability of developing robust models for specific scenarios and/or architectures in the context of HTR.
-The “Flip” method with cropped patches specifically stands out, making it the go-to strategy for those seeking to achieve the most robust performance.
-Figure 4: Comparison of occurrences of the proposed methods in the semi-supervised scenario in which they outperform the state-of-the-art ChaCo method [81] for any data collection and architectural framework (TD or CTC).
-The figures are provided as a ratio with respect to the total number of experimental scenarios considered.
-The dashed line indicates the chance of our methods to surpass the ChaCo strategy.
+(a) Rotation
+(b) Flip
+(c) Jigsaw
+(d) Sorting
+Figure 4: Heat map highlighting the regions targeted by the encoder pre-trained with the different spatial context-based SSL methods.
+At last, although many of our proposals reflect robust behavior that equal on the state of the art, the “Flip” method is positioned as the best-performing alternative in general terms.
+This method not only surpasses other spatial context-based techniques but also outperforms prior SSL methods in the semi-supervised scenario.
+In particular, “Flip" exhibits clear superiority over the “Rotation" strategy when using CTC decoding.
+We believe that this advantage arises from Flip’s preservation of the horizontal orientation of text, which aligns well with CTC’s horizontal decoding of feature representations.
+This reinforces the value of developing and implementing strategies that are directly aligned to the unique challenges and characteristics of HTR task.
+6.2 Computational architectural requirements
+While CRNN architectures with CTC decoding have dominated the state of the art for years, transformer architectures have recently demonstrated exceptional effectiveness in HTR tasks mainly due to their low inductive bias and language model ability.
+Nevertheless, this advantage comes at the expense of increased computational complexity and resource demands [23].
+Given these trade-offs, it is essential to assess the efficiency and practicality of the chosen architecture.
+Therefore, this section analyzes the computational requirements of the architectures considered in the work and compares them with the state-of-the-art SSL methods.
+Table 4: Table showing the input size, GFLOPs, and number of parameters for each method.
+Architecture Input size GFLOPs Params
+CTC (ours) [64,] 20.4 48.0M
+TD (ours) [64,] 20.7 50.7M
+[9] CTC [32, 100] 9.7 48.5M
+[9] Att. [32, 100] 24.5 49.5M
+Text-DIAE [64, 256] 27.6 68.8M
+Table 4 presents the Floating Point Operations (FLOPs) and the number of parameters for the architectures used by the different methods for which we were able to replicate the architecture.
+Focusing first on the architectures employed in this work, as expected, the one incorporating a TD decoder exhibits higher FLOP consumption.
+However, since the decoder consists of only one layer, the additional computational cost remains minimal—just 0.3 GFLOPs.
+This makes it a highly favorable trade-off, offering a modest increase in resource usage and parameters while delivering a substantial improvement in performance.
+In contrast, Text-DIAE, a Transformer-based method, emerges as the most resource-intensive architecture among the evaluated but it does not deliver the best results as shown in Table 3(b), underscoring the fact that simpler architectures and methods can surpass these more complex approaches in performance.
+This highlights the importance of balancing architectural complexity with efficiency and effectiveness in HTR systems.
+On the other hand, when using CTC, our models exhibit more FLOP consumption compared to other CRNN-based methods (note that ChaCo, SeqCLR and CMT-Co share the architecture of [9]) despite having a similar number of parameters.
+This is because the input image size in our case is, typically, larger.
+However, this issue is notable mitigated by the inherent computational simplicity of the methods, as we will discuss in next section.
+6.3 Strengths and limitations
+In recent years, contrastive learning or dual-branch architectures has significantly advanced capabilities in image classification.
+However, its direct application to HTR is not straightforward.
+For example, aggressive data augmentation can cause misalignment between positive and negative pairs, leading to suboptimal representations [1].
+In contrast, the spatial context-based SSL methods explored in this work avoid such issues.
+By focusing solely on stroke patterns within the processed images, these methods can leverage far more aggressive data augmentations, such as strong geometric transformations or croppings.
+This approach enables the network to encounter a wider variety of data, fostering a deeper understanding of their intrinsic characteristics.
+Additionally, the simplicity of the presented pretext tasks—e.g., classifying an image crop in terms of its rotation angle—reduces computational training costs compared to dual-branch architectures or generative approaches while achieving similar performance rates.
+Concerning the limitations, it has also been proven that dual-branch architectures or generative methods show a higher abstraction capacity in the representations by encapsulating more semantic information.
+In this regard, since the presented proposals exclusively focus on the visual information extracted from the stroke and font style, its applicability to other off-the-shelf tasks—e.g., text segmentation, writer retrieval or image enhancement—remains uncertain.
+Future works could explore the generalizability of the different SSL methods.
 7 Conclusions
 In this paper, we study spatial context-based SSL methods originally devised for classification tasks and adapt them to the HTR field.
-We propose two possible adaptation frameworks for this purpose: (i) adjusting the data in hand in order to use the original SSL methods (input adaptation), and (ii) adapting these SSL approaches to the sequential nature of HTR (method adaptation).
-The results obtained when considering three benchmark HTR datasets—IAM, CVL and George Washington—show that these proposals not only improve the recognition rates of the model when not considering any pre-training, but also consistently outperform the current state-of-the-art SSL methods for HTR.
-Moreover, it should be noted that one of our proposals that was specifically designed for HTR (flip) consistently obtains the overall best results in every experimental scenario.
-Note also that our results are aligned with those of previous SSL literature on text recognition [49], which argues that there is a close relationship between HTR and spatial context owing to stroke and font style information.
-The objective of this study is to provide practical guidelines in order to take advantage of SSL techniques in the context of HTR, thus combating the problem of insufficient data in this field.
-Appendix A Data augmentation
-We detail the data augmentation techniques applied in our study, leveraging the imgaug augmentation package [35].
-Our selection is informed by seminal contributions in the field [3, 81], leading to the adoption of the subsequent alterations:
+We propose two possible adaptation frameworks for this purpose: (i) adjusting the data in order to use the original SSL methods (input adaptation), and (ii) adapting these SSL approaches to the sequential nature of HTR (method adaptation).
+The results obtained when considering three benchmark HTR datasets—IAM, CVL, and RIMES—show that these proposals achieve state-of-the-art performance in various scenarios, particularly in semi-supervised settings.
+Moreover, among the proposals, it should be noted that one specifically designed for HTR (“Flip”) seems to be specially effective.
+Also note that our results are aligned with those of previous SSL literature on text recognition [25], which argues that there is a close relationship between HTR and spatial context owing to stroke and font style information.
+Then, the final objective of this study is to provide practical guidelines to take advantage of a family of SSL techniques in HTR context, thus encouraging the study and application of different SSL frameworks in this field.
+Appendix A Experimental set-up
+In this section we expand the experimental configuration considered in the work.
+More precisely, we detail the data augmentation processes, the visual encoder architecture, and the partitioning of the RIMES dataset used.
+A.1 Data augmentation
+We leverage hard data augmentation techniques enabled by the SSL methods employed, complemented by additional augmentation strategies inspired by seminal works in the field [1, 41].
+This approach results in the adoption of the following transformations by using the imgaug augmentation package [19]:
+1aug = iaa.Sequential([
+2 iaa.SomeOf((1, 8), [
+3 # Geometric
+4 iaa.ScaleX((0.75, 1.0), mode=’constant’, cval=255),
+5 iaa.ScaleY((0.75, 1.25), mode=’constant’, cval=255),
+6 iaa.TranslateX(percent=(-0.1, 0.1), mode=’constant’, cval=255),
+7 iaa.TranslateY(percent=(-0.15, 0.15), mode=’constant’, cval=255),
+8 iaa.Rotate((-5., 5.), mode=’constant’, cval=255),
+9 iaa.ShearX((-25, 25), mode=’constant’, cval=255),
+10 iaa.PiecewiseAffine(scale=(0.02, 0.04)),
+11 # Color
+12 iaa.Sharpen(alpha=(0.1, 0.5), lightness=(0.75, 1.25)),
+13 iaa.LinearContrast((0.5, 1.5)),
+14 # Blur
+15 iaa.MotionBlur(k=(5, 7)),
+16 iaa.GaussianBlur((1.0, 2.0)),
+17 # Noise
+18 iaa.AdditiveGaussianNoise(scale=(1, 25))
+19 ], random_order=False)
+20])
+21# Morphological
+22num = random.uniform(0, 1)
+23if num < 0.33:
+24 image = erode(image)
+25elif random.uniform(0, 1) < 0.66:
+26 image = dilate(image)
+27image = aug(image=image)
 Listing 1: Data augmentation used in the SSL stage.
-iaa.Sequential([
-iaa.SomeOf((1, 5), [
-iaa.LinearContrast((0.5, 1.0)),
-iaa.GaussianBlur((0.5, 1.5)),
-iaa.Sharpen(alpha=(0.0, 0.5), lightness=(0.0, 0.5)),
-iaa.PiecewiseAffine(scale=(0.02, 0.03), mode=’edge’),
-iaa.PerspectiveTransform(scale=(0.01, 0.02)),
-], random_order=True)]
 Examples of the transformations generated by the previous code are shown in Fig. 5.
 Figure 5: Examples of data augmentations.
-Furthermore, drawing on established practices in the literature, we also employ a soft data augmentation strategy during the supervised stage.
-This includes subtle cropping, linear contrast adjustments, and Gaussian blur.
-Appendix B Detailed results
-Owing to the extensive experimentation carried out in this paper, this section offers a complete compilation of results for each method with respect to the specific architectures employed and the datasets used for pre-training and evaluation.
-In Tab. B we report the quality results in terms of the CER and WAcc metrics.
-In Tab.
-B we provide the semi-supervised results in terms of WAcc and Single Edit Distance (ED1) metrics.
-ED1 expresses word-level accuracy up to a single edit operation.
-Note that, the exclusion of the latter metric in the manuscript was primarily due to concerns regarding clarity.
-Nevertheless, in this section, we opted to include it to enhance the comprehensiveness of the results presented.
-It is worth highlighting that the incorporation of this metric does not deviate from the conclusions drawn in the paper; instead, it serves to reinforce and bolster them.
-CER WAcc
-IAM CVL G.W. IAM CVL G.W.
-Geometric
-Rotation 27.4 19.7 35.1 53.6 56.8 25.5
-Flip 23.8 20.5 28.8 56.1 57.3 43.8
-[…]
-Jigsaw 52.4 28.2 31.3 31.6 46.0 42.2
-Sorting 79.7 53.0 31.4 19.4 12.3 43.2
-ChaCo 25.2 23.7 84.7 48.3 47.0 04.7
-(a) Representation quality under CTC decoder.
-CER WAcc
-IAM CVL G.W. IAM CVL G.W.
-Geometric
-Rotation 19.0 11.9 22.6 64.6 79.2 59.8
-Flip 17.4 15.3 19.3 64.7 75.1 61.1
-[…]
-Jigsaw 49.1 19.6 38.3 35.7 70.5 44.9
-Sorting 72.1 58.8 64.0 8.6 24.7 19.8
-ChaCo 12.2 11.9 69.2 71.6 80.2 10.0
-(b) Representation quality under TD decoder.
-(c) Representation quality. The frozen part is only the visual encoder pre-trained in the SSL method. Bold values indicate the best results under the same conditions of pre-training and test.
-WAcc ED1
-IAM CVL G.W. IAM CVL G.W.
-5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100%
-Baseline
-No pretrain 26.5 42.2 70.4 07.5 14.2 69.7 13.5 16.0 78.5 42.1 64.9 88.1 17.4 25.4 85.6 26.3 44.6 92.3
-ChaCo 30.0 46.3 64.6 13.1 13.2 67.4 15.7 21.2 72.6 47.7 70.2 83.9 34.6 40.7 84.0 28.7 34.9 89.6
-[…]
-Rotation 42.6 48.9 67.7 10.0 16.4 70.1 17.0 36.0 82.4 67.7 74.4 88.2 27.2 38.8 85.9 32.9 56.9 94.0
-Flip 42.1 50.6 72.1 17.1 35.0 71.1 18.8 32.5 79.1 68.4 75.6 89.4 38.7 61.1 86.8 35.9 54.3 93.2
-Jigsaw 42.4 51.4 67.7 05.3 12.8 55.4 12.9 19.5 72.9 66.8 75.2 86.7 18.1 31.2 77.3 26.4 43.9 89.6
-Sorting 23.9 41.6 61.1 09.3 14.7 70.0 13.8 08.9 76.6 38.1 63.9 81.1 19.3 26.2 84.7 24.1 33.1 91.4
-(d) Semi-supervised results under CTC decoder.
-WAcc ED1
-IAM CVL G.W. IAM CVL G.W.
-5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100%
-Baseline
-No pretrain 10.3 50.3 74.8 05.9 14.2 68.7 06.3 02.1 71.3 21.7 63.6 87.6 07.9 25.4 85.6 09.2 02.4 82.8
-ChaCo 47.3 53.8 75.1 21.2 38.0 83.4 05.9 07.0 73.2 66.6 74.0 88.2 27.1 44.9 87.9 06.6 08.2 83.1
-[…]
-Rotation 47.0 54.5 74.5 21.5 51.8 85.5 03.6 06.9 78.5 65.6 74.8 87.2 28.3 61.4 90.6 07.7 08.7 89.5
-Flip 43.8 54.3 74.6 25.2 53.1 86.1 06.0 11.0 78.0 60.2 72.0 88.1 29.7 62.8 90.7 08.9 13.5 90.3
-Jigsaw 43.5 51.7 74.3 21.5 38.1 84.8 14.6 20.2 77.1 60.5 69.1 87.1 27.8 44.6 88.5 18.9 26.9 88.7
-Sorting 25.2 39.8 75.2 05.4 22.0 83.6 11.9 14.2 76.1 34.5 49.4 87.2 07.0 26.5 87.0 15.4 17.2 86.6
-(e) Semi-supervised results under TD decoder.
-(f) Semi-supervised results.
-The models pre-trained with the SSL methods are fine-tuned with 5%, 10% and 100% of the labelled data.
-Bold values indicate the best results under the same conditions of pre-training and test.
+Furthermore, drawing on established practices in the literature, we also employ the data augmentation strategy during the supervised stage.
+This includes the previous data augmentations excluding hard geometric augmentations.
+A.2 Visual encoder architectural details
+In this subsection, Table 5 details the precise configuration of the encoder architecture contemplated in the work.
+Note that, as commented in the manuscript, this scheme stands for a ResNet-29 architecture [9] with slight modifications.
+Table 5: Architectural details of the neural encoder considered. Block refers to a block of convolutional layers.
+Layer Configuraton
+Kernel, Channels Stride Padding
+Convolution $3\times 3$, 32 (1, 1) same
+Convolution $3\times 3$, 64 (1, 1) same
+Max pooling $2\times 2$, - (2, 2) (0, 0)
+Block $\begin{bmatrix}3\times 3,&128\\ 3\times 3,&128\end{bmatrix}$$\times 1$ (1, 1) same
+Max pooling $2\times 2$, - (2, 2) (0, 0)
+Block $\begin{bmatrix}3\times 3,&256\\ 3\times 3,&256\end{bmatrix}$$\times 2$ (1, 1) same
+Max pooling $2\times 2$, - (2, 2) (0, 0)
+Block $\begin{bmatrix}3\times 3,&512\\ 3\times 3,&512\end{bmatrix}$$\times 5$ (1, 1) same
+Convolution $3\times 3$, 512 (1, 1) same
+Max pooling $2\times 2$, - (2, 2) (0, 0)
+Block $\begin{bmatrix}3\times 3,&512\\ 3\times 3,&512\end{bmatrix}$$\times 3$ (1, 1) same
+Max pooling $2\times 2$, - (2, 2) (0, 0)
+Convolution $3\times 3$, 512 (1, 1) same
+A.3 RIMES partition details
+In this work, we utilized the latest version of the RIMES word dataset [15].
+This dataset is organized into 57 folders, each referred to as a block, containing each one 100 letters represented as word-level images.
+To our best knowledge, there is no official partitioning scheme for this version of the dataset, nor are there prior studies that use it.
+Consequently, we have created arbitrary partitions for this work whose details are now provided to ensure reproducibility and facilitate future comparisons:
+Table 6: Table showing the partitions employed in RIMES dataset
+Partition Blocks Samples
+Train 1-34 139490
+Validation 35-39 20217
+Test 40-57 68651

### v3

@@
 (b) We propose two new approaches adapted to the specific features of HTR: (i) “Flip” as a geometric transformation; (ii) “Sorting” as a puzzle transformation.
-This work was motivated by the relevance of spatial information in HTR tasks [25] together with the fact that, to our best knowledge, it has not been previously studied, and focuses on the adaptation of a reference set of such approaches to the HTR field.
+This work was motivated by the relevance of spatial information in HTR tasks [11, 26] together with the fact that, to our best knowledge, it has not been previously studied, and focuses on the adaptation of a reference set of such approaches to the HTR field.
 More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig.
@@
 Despite being largely explored in classification frameworks, spatial context-based SSL is rarely exploited in image-based text recognition tasks.
-Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—suggest that, if adequately adapted, text recognition models should benefit from pre-training methods of this nature.
+Nevertheless, the competitive results that are obtained by employing these strategies for classification, together with the strong relationship between spatial information and HTR—owing to stroke and font style [11]—suggest that, if adequately adapted, text recognition models should benefit from pre-training methods of this nature.
+Then, while previous efforts on SSL for HTR have focused on contrastive or image generation frameworks, this work explores the spatial context-based paradigm.
 Note that, our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
@@
 Therefore, this section analyzes the computational requirements of the architectures considered in the work and compares them with the state-of-the-art SSL methods.
-Table 4: Table showing the input size, GFLOPs, and number of parameters for each method.
+Table 4: Table showing the input size, GFLOPs, and number of parameters for each method. Symbol * denotes that the number of columns is dynamically adjusted to preserve the image aspect ratio.
 Architecture Input size GFLOPs Params
-CTC (ours) [64,] 20.4 48.0M
-TD (ours) [64,] 20.7 50.7M
+CTC (ours) [64,*] 20.4 48.0M
+TD (ours) [64,*] 20.7 50.7M
 [9] CTC [32, 100] 9.7 48.5M
@@
 Concerning the limitations, it has also been proven that dual-branch architectures or generative methods show a higher abstraction capacity in the representations by encapsulating more semantic information.
-In this regard, since the presented proposals exclusively focus on the visual information extracted from the stroke and font style, its applicability to other off-the-shelf tasks—e.g., text segmentation, writer retrieval or image enhancement—remains uncertain.
+In this regard, since the presented proposals exclusively focus on the visual information extracted from the stroke and font style, its applicability to other off-the-shelf tasks—e.g., text segmentation, writer retrieval, image enhancement, multilingual or historical manuscripts—remains uncertain.
 Future works could explore the generalizability of the different SSL methods.